# Compact Kähler manifolds with quasi-positive holomorphic sectional curvature

### v3

@@
Abstract. In this paper, we establish a “pseudo-effective” version of the holonomy principle for compact Kähler manifolds with nonnegative holomorphic sectional curvature. As applications, we prove that if a compact complex manifold $M$ admits a Kähler metric with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), $M$ is projective and rationally connected. This answers a problem raised by Yang and Matsumura and extends Yau’s conjecture. We also get the structure theorem of a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature. Furthermore, we show that a compact Kähler manifold admits a Hermitian metric with positive real bisectional curvature is a projective and rationally connected manifold.
@@ 1. Introduction
A projective manifold $M$ is called rationally connected if any two points on $M$ can be connected by some rational curve. The rational connectedness is an important and fundamental concept in algebraic geometry, and many people have given the criteria for it ([8, 17, 31, 23, 7, 22]). It is also interesting to give geometric interpretations for rational connectedness. Let $\omega$ be a Kähler metric on $M$, the holomorphic sectional curvature of $\omega$ along the nonzero $V\in T^{1,0}M$ is defined by
@@
where $R$ is the Riemannian curvature tensor of $\omega$. S.-T. Yau proposed the following well-known conjecture in his problem section.
Conjecture 1.1 ([39, Problem 47]).
@@
By studying MRC fibration, Heier and Wong ([18]) confirmed Yau’s conjecture for projective manifolds with quasi-positive holomorphic sectional curvature. Then Yang ([36, 37]) proved that a Kähler manifold $M$ with positive holomorphic sectional curvature satisfies $H^{p,0}_{\bar{\partial}}(M)=H^{0,2}_{\bar{\partial}}(M)=0$ by introducing the notion of RC-positivity and affirmly confirmed Yau’s conjecture, but it’s seems difficult to apply his method to the case of quasi-positive holomorphic sectional curvature. There are some other interesting curvature conditions toward rational connectedness. L.Ni ([28]) introduced the concept of $k$-Ricci curvature for compact Kähler manifold and generalized Yang’s method to show that a compact Kähler manifold with positive $k$-Ricci curvature is projective and rationally connected. Recently, Li, Zhang and the second author ([24, Theorem 1.14]) established the equivalence between mean curvature (quasi-)positivity and rational connectedness for compact Kähler manifolds.
In [25, 26], Matsumura established the structure theorem for a projective manifold $X$ with nonnegative holomorphic sectional curvature by studying the structure of MRC fibration: the universal cover of $X$ splits isometrically as $F\times\mathbb{C}^{n-k}$ for some rationally connected manifold $F$. As a corollary, he generalized Yau’s conjecture for projective manifolds ([26, Theorem 1.2]): if a projective manifold $M$ admits a Kähler metric $\omega$ with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), then $M$ is rationally connected. Following the definition in [20], a tangent vector $V\in T_{x_{0}}^{1,0}M$ is said to be truly flat at $x_{0}$ with respect to $\omega$ if
$R(V,\overline{X},Y,\overline{Z})=0$
for all $X,Y,Z\in T_{x_{0}}^{1,0}M$. Motivated by these results, Matsumura and Yang proposed the following natural generalization.
Conjecture 1.2 ([37, Conjecture 1.9], [27, Problem 3.3]).
Let $M$ be a compact Kähler manifold. If $M$ admits a Kähler metric (or more generally a Hermitian metric) with nonnegative holomorphic sectional curvature and there has no nonzero truly flat tangent vector at some point (or the holomorphic sectional curvature is quasi-positive), then $M$ is projective and rationally connected.
More generally, Matsumura proposed the following extension and suggested to investigate Albanese maps instead of MRC fibrations for compact Kähler manifolds. We refer the readers to [27, Section 3], [26, Section 5] for more detailed discussion.
Problem 1.3 ([27, Problem 3.5]).
@@
In this paper, we establish the following “pseudo-effective” version of the holonomy principle for nonnegative holomorphic sectional curvature (cf. [8, Theorem 1.5]), which is the first main result of our paper.
@@
Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature. Assume that there exists an invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ that is pseudo-effective as a line bundle, then the following statements holds:
@@
    There exists some nonzero truly flat tangent vectors at any point of $M$ determined by $\mathcal{F}$.
@@
    $\mathcal{F}$ is flat as a line bundle and is invariant under parallel transport by the connection $D$ of $\Lambda^{p,0}M$ induced by Chern connection of $(T^{1,0}M,\omega)$; in fact, the local holonomy $H$ of $(T^{1,0}M,\omega)$ acts trivally on $\mathcal{F}$.
The proof of Theorem 1.4 is based on an integral inequality associated with $\mathcal{F}$ (Proposition 3.1). As a direct application of Theorem 1.4 (1) and Campana-Demailly-Peternell’s criterion for rational connectedness, we get the second main result of our paper, which confirms Conjecture 1.2 for the Kähler case and extends Yau’s conjecture to the quasi-positive case. Our approach is independent of those of Heier-Wong ([18]) and Matsumura ([26]) for projective manifolds.
@@
Statement (2) of Theorem 1.4 is same as the statement of [8, Theorem 1.5] for nonnegative Ricci curvature, but the proof is more involved. Based on this statement, Campana-Demailly-Peternell’s criterion for rational connectedness (Lemma 2.5), Cheeger-Gromoll’s splitting theorem ([9]), De Rham’s splitting theorem ([15]) and Berger’s classification ([2]) for irreducible holonomy, Campana, Demailly and Peternell established the structure theorem of a compact Kähler manifold $M$ with nonnegative Ricci curvature ([8, Theorem 1.4]): the universal cover $\widetilde{M}$ of $M$ can splits isometrically as $\mathbb{C}^{q}\times\prod Y_{j}\times\prod S_{k}\times\prod Z_{l}$, where $Y_{j}$, $S_{k}$ and $Z_{l}$ are compact simply connected Kähler manifolds of respective dimensions $n_{j}$, $n^{\prime}_{k}$, $n^{\prime\prime}_{l}$ with irreducible holonomy, $Y_{j}$ being Calabi-Yau manifolds (holonomy $SU(n_{j})$), $S_{k}$ holomorphic symplectic manifolds (holonomy $Sp(n^{\prime}_{k}/2)$), and $Z_{l}$ rationally connected manifolds with $K_{Z_{l}}^{-1}$ semipositive (holonomy $U(n^{\prime\prime}_{l})$, unless $Z_{l}$ is Hermitian symmetric of compact type).
Statement (1) of Theorem 1.4 ensures that factors of irreducible holonomy $H=SU(n_{j})$ or $Sp(\frac{n_{k}^{\prime}}{2})$ can be excluded for a compact Kähler manifold with nonnegative holomorphic sectional curvature.
@@
Let $(M,\omega)$ be a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature. Then $M$ admits a holomorphic and isometric splitting $(M,\omega)\simeq\prod(M_{l},\omega_{l})$ where $M_{l}$ are rationally connected manifolds with nonnegative holomorphic sectional curvature and irreducible holonomy (holonomy $U(n_{l})$ unless $M_{l}$ is Hermitian symmetric of compact type).
If the following splitting theorem for nonnegative holomorphic sectional curvature established, Problem 1.3 will be resolved.
Question 1.8.
@@
For the Hermitian category, Yang ([36, Corollary 1.10]) proved that a compact Kähler surface admits a Hermitian metric with positive holomorphic sectional curvature is projective and rationally connected. In [38], Yang-Zheng introduced the concept of real bisectional curvature as in the following way and generalized Wu-Yau’s Theorem ([19, 34, 35, 16]) to the Hermitian case. The underlying reason for the extension is that the Schwarz lemma works the same when the target metric is only Hermitian but has nonpositive real bisectional curvature on a Kähler background. Also, they proposed the following generalization of Yau’s Conjecture 1.1 for the Hermitian case ([38, Conjecture 1.6 (b)]).
@@
We say that a Hermitian manifold $(M,h)$ has positive real bisectional curvature denoted by $B_{h}>0$ if for any $x\in M$, any unitary frame $e=\{e_{1},\cdots,e_{n}\}$ at $x$ and any nonnegative constants $a=\{a_{1},\cdots,a_{n}\}$ with $|a|^{2}=a_{1}^{2}+\cdots+a_{n}^{2}>0$, it holds that
@@
where $R$ is the chern curvature tensor of $h$. The term $B_{h}\geq 0,\ B_{h}<0$ or $B_{h}\leq 0$ are defined similarly. This curvature condition is the same as positive holomorphic sectional curvature when the metric $h$ is Kähler and slightly stronger in the non-Kähler case (cf. [38]). Under the entra assumption that $M$ is Kählerian, Conjcture 1.9 is just a weaker version of Conjecture 1.2. In this special case, Tang ([33, Theorem 1.4]) confirmed the Conjecture 1.9 for $\dim M=3$ and proved that $M$ is projective for $\dim M\geq 1$. Since the curvature term of Bochner-type formula in the proof of Proposition 3.1 can be controlled by real bisectional curvature, we confirm Conjecture 1.9 under the additional assumption that $M$ is Kählerian.
@@
This paper is organized as follows. In section 2, we recall some basic definitions and facts. In section 3, we give a proof of Theorem 1.4 by establishing an integral inequality. In section 4 , we give a proof of Theorem 1.5 and Theorem 1.7. In section 5, we give a proof of Theorem 1.10.
@@ 2. Preliminaries
For reader’s convenience, we gather some basic definitions and useful facts known to experts, which will be referenced later in the text.
@@ 2.1. Pseudo-effective line bundle and plurisubharmonic function
Let $\Omega$ be an open subset of $\mathbb{C}^{n}$. We have the following approxiamation properties for plurisubharmonic functions (see e.g. [13, Chapter 1 and Chapter 3]).
@@
([3, Theorem 2.1]) Let $\{v_{j}^{1}\},\cdots,\{v_{j}^{k}\}$ be a decreasing sequences of functions in $Psh(\Omega)\cap L^{\infty}_{loc}(\Omega)$ and assume that for all $z\in\Omega$,
@@ 2.2. Some criterions for rational connectedness
Based on the uniruledness criterion estabished in ([6]), Campana-Demailly-Peternell establish the following criterions for rational connectedness.
Lemma 2.5 ([8, Theorem 1.1]).
@@
    For every invertible subsheaf $\mathcal{F}\subset\mathcal{O}(T^{1,0}M^{*})^{\otimes p})$, $p\geq 1$, $\mathcal{F}$ is not pseudo-effective.
  4. (d)
@@ 2.3. Some properties of curvature tensor
(Royden’s observation, [32, P552]) Let $\xi_{1},\cdots,\xi_{k}$ be $k$ orthonormal tangent vectors. If $S(\xi,\bar{\eta},\zeta,\bar{\omega})$ is a symmetric bihermitian form on some complex vector space $V$ (i.e., $S(\xi,\bar{\eta},\zeta,\omega)=S(\zeta,\bar{\eta},\xi,\bar{\omega})$ and $S(\xi,\bar{\omega},\zeta,\bar{\eta})$), such that for all $\xi\in V$, we have
@@
([37, Lemma 1.6]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$. Let $e_{1}\in T_{x_{0}}^{1,0}M$ be a unit vector which minimizes the holomorphic sectional curvature $H$ of $\omega$ at $x_{0}$, then
@@
([20, Lemma 2.1]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$. Assume that the holomorphic sectional curvature $H$ of $\omega$ is nonnegative at $x_{0}$. If a nonzero tangent vector $V\in T^{1,0}_{x_{0}}M$ such that $R(V,\overline{V},W,\overline{W})=0$ for any $W\in T^{1,0}_{x_{0}}(M)$, then
@@
3. Proof of Theorem 1.4
In this subsection, we prove Theorem 1.4. The proof is divided into two parts: establishing an integral inequality and performing a pointwise computation. The method is based on integration by parts, Bochner-type formula and algebraic properties of holomorphic sectional curvature.

### v4

# On the structure of compact Kähler manifolds with nonnegative holomorphic sectional curvature

@@
Abstract. In this paper, we establish a “pseudo-effective” version of the holonomy principle for compact Kähler manifolds with nonnegative holomorphic sectional curvature. As applications, we prove that if a compact complex manifold $M$ admits a Kähler metric $\omega$ with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), then $M$ must be projective and rationally connected. This result answers a problem raised by Yang and Matsumura and extends the Yau’s conjecture. We also prove that a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature is projective and rationally connected. Additionally, we classify non-projective compact Kähler 3-dimensional manifolds. Furthermore, we show that a compact Kähler manifold admits a Hermitian metric with positive real bisectional curvature is a projective and rationally connected manifold.
@@ 1. Introduction
A projective manifold $M$ is called rationally connected if any two points on $M$ can be connected by some rational curve. The rational connectedness is an important and fundamental concept in algebraic geometry, and many people have given the criteria for it ([9, 19, 37, 27, 8, 26]). It is also important to provide geometric interpretations for rational connectedness. Let $\omega$ be a Kähler metric on $M$, the holomorphic sectional curvature of $\omega$ along a nonzero $V\in T^{1,0}M$ is defined by
@@
where $R$ is the Riemannian curvature tensor of $\omega$. S.-T. Yau proposed the following well-known conjecture in his problem section:
Conjecture 1.1 ([45, Problem 47]).
@@
Heier and Wong ([21]) confirmed Yau’s conjecture for projective manifolds with quasi-positive holomorphic sectional curvature by studying MRC fibration. By introducing the notion of RC-positivity and studying vanishing theorems, Yang ([42, 43]) showed that a Kähler manifold $M$ with positive holomorphic sectional curvature is projective and rationally connected, then affirmly confirmed Yau’s conjecture.
There are some other interesting curvature conditions related to rational connectedness (see, for instance, [34, 36, 42, 43]). Recently, Li, Zhang and the second author ([29, Theorem 1.14]) established the equivalence between mean curvature quasi-positivity of the tangent bundle and rational connectedness for compact Kähler manifolds.
In [30, 31], Matsumura established a structure theorem for projective manifolds $X$ with nonnegative holomorphic sectional curvature by studying the structure of MRC fibration. Both the methods of Heier-Wong and Matsumura rely on the uniruledness criterion for projective manifolds ([7]). According to Matsumura’s structure theorem, the universal cover of $X$ splits isometrically and holomorphically as $F\times\mathbb{C}^{n-k}$, where $F$ is a rationally connected manifold. This result generalizes the structure theorem established by Howard, Smith and Wu ([24]) and Mok ([33]) for nonnegative holomorphic bisectional curvature. As a corollary, Matsumura generalized Yau’s conjecture for projective manifolds ([31, Theorem 1.2]): if a projective manifold $M$ admits a Kähler metric $\omega$ with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), then $M$ is rationally connected. According to [23], a tangent vector $V\in T_{x_{0}}^{1,0}M$ is said to be truly flat at $x_{0}$ with respect to $\omega$ if
$R(V,\overline{X},Y,\overline{Z})=0,\text{\ for all $X,Y,Z\in T^{1,0}M$}.$
Motivated by these results, Matsumura and Yang proposed the following natural generalization.
Conjecture 1.2 ([43, Conjecture 1.9], [32, Problem 3.3]).
Let $M$ be a compact Kähler manifold. If $M$ admits a Kähler metric (or more generally a Hermitian metric) with nonnegative holomorphic sectional curvature and there has no nonzero truly flat tangent vector at some point, then $M$ is projective and rationally connected.
More generally, Matsumura proposed the following extension and suggested investigating Albanese maps instead of MRC fibrations for compact Kähler manifolds, due to the lack of a uniruledness criterion in the Kähler case. We refer the reader to [32, Section 3] and [31, Section 5] for a more detailed discussion.
Problem 1.3 ([32, Problem 3.5]).
@@
Towards Conjecture 1.2 and Problem 1.3, we establish a “pseudo-effective” version of the holonomy principle for nonnegative holomorphic sectional curvature (c.f. [9, Theorem 1.5] for nonnegative Ricci curvature). This is the first main result of our paper.
@@
Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature. Assume there exists an invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ that is pseudo-effective as a line bundle. Then the following statements holds:
@@
    $\mathcal{F}$ is flat as a line bundle and invariant under parallel transport by the connection $D$ of $\Lambda^{p,0}M$ induced by Chern connection of $(T^{1,0}M,\omega)$. In fact, the local holonomy $H$ of $(T^{1,0}M,\omega)$ acts trivally on $\mathcal{F}$.
@@
    $\mathcal{F}$ determines a parallel and self-adjoint linear tranformation $S:T^{1,0}M\rightarrow T^{1,0}M$, which induces a decomposition $V+W$ of $T^{1,0}M$ where $V$ and $W$ consist of eigenvectors corresponding to nonzero and zero eigenvalues of $S$, respectively; in fact, $\forall p\in M$, every $v\in V_{p}$ is truly flat and $\dim V_{p}=\operatorname{rank}(S)>0$.
The core of the proof of Theorem 1.4 is an integral inequality associated with $\mathcal{F}$ (Proposition 3.1). As a direct application of Theorem 1.4 and the Campana-Demailly-Peternell’s criterion for rational connectedness (2.5), we obtain the second main result of our paper, which confirms Conjecture 1.2 for Kähler manifolds and extends Yau’s conjecture to the quasi-positive case. Our approach is independent of those by Heier-Wong ([21]) and Matsumura ([31]) for projective manifolds.
@@
Statement (1) of Theorem 1.4 parallels the statement of [9, Theorem 1.5] for nonnegative Ricci curvature, but the proof is more involved. Based on this statement, Campana-Demailly-Peternell’s criterion for rational connectedness (Lemma 2.5), Cheeger-Gromoll’s splitting theorem ([10]), De Rham’s splitting theorem ([17]) and Berger’s classification ([3]) for irreducible holonomy, Campana, Demailly and Peternell established the structure theorem of a compact Kähler manifold $M$ with nonnegative Ricci curvature ([9, Theorem 1.4]): the universal cover $\widetilde{M}$ of $M$ can splits isometrically as $\mathbb{C}^{q}\times\prod Y_{j}\times\prod S_{k}\times\prod Z_{l}$, where $Y_{j}$, $S_{k}$ and $Z_{l}$ are compact simply connected Kähler manifolds of respective dimensions $n_{j}$, $n^{\prime}_{k}$, $n^{\prime\prime}_{l}$ with irreducible holonomy. Specifically, $Y_{j}$ are Calabi-Yau manifolds (holonomy $SU(n_{j})$), $S_{k}$ are holomorphic symplectic manifolds (holonomy $Sp(n^{\prime}_{k}/2)$), and $Z_{l}$ are rationally connected manifolds with $K_{Z_{l}}^{-1}$ semipositive (holonomy $U(n^{\prime\prime}_{l})$, unless $Z_{l}$ is Hermitian symmetric of compact type).
Comparing with [9, Theorem 1.4] regarding nonnegative Ricci curvature, Ricci-flat factors can be excluded for a compact Kähler manifold with nonnegative holomorphic sectional curvature. As another application of Theorem 1.4, we affirm Problem 1.3 for the simply connected case, which constitutes the third main result of our paper.
@@
If $(M,\omega)$ is a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature, then $M$ is projective and rationally connected.
According to Theorem 1.7, Problem 1.3 is equivalent to the following splitting theorem for nonnegative holomorphic sectional curvature, which is an analogue of Cheeger-Gromoll’s splitting theorem ([10]) for nonnegative Ricci curvature.
Problem 1.8.
@@
Matsumura has affirmed Problem 1.3 for compact Kähler surfaces ([30, Corollary 1.5]). To support Problem 1.3 and 1.8 in dimension 3, we provide the following classification of non-projective Kähler 3-dimensional manifolds with nonnegative holomorphic sectional curvature by considering Albanese maps. This classification can also serve as a concrete example.
Theorem 1.9.
Let $(M,\omega)$ be a non-projective compact Kähler 3-dimensional manifold with nonnegative holomorphic sectional curvature. Then $M$ is either flat or a $\mathbb{P}^{1}$-bundle over a torus $\mathbb{T}^{2}$. In the latter case, the universal cover of $M$ is $\mathbb{P}^{1}\times\mathbb{C}^{2}$ endowed with the product of a Kähler metric with quasi-positive holomorphic sectional curvature on $\mathbb{P}^{1}$ and the flat metric on $\mathbb{C}^{2}$.
Remark 1.10.
From Proposition 4.2 that was used in the proof of Theorem 1.9, it is likely that there exists a parallel result similar to [16, Proposition 3.10] for compact Kähler manifolds with nonnegative holomorphic sectional curvature. If this is the case, then Problem 1.3 could be resolved by considering Albanese maps as discussed in [16, Section 3]. However, we currently do not have any information about the fundamental group of such manifolds. Therefore, from this perspective, it might be more necessary and direct to consider Problem 1.8.
Finally, we consider the Hermitian case. Yang ([42, Corollary 1.10]) proved that a compact Kähler surface with a Hermitian metric of positive holomorphic sectional curvature is projective and rationally connected. It remains difficult to consider the Hermitian case in dimension $\geq$ 3. In [44], Yang and Zheng introduced the concept of real bisectional curvature and generalized Wu-Yau’s Theorem ([22, 40, 41, 18]) to the Hermitian case. This concept is also referred as ”positive curvature operator” by Lee and Streets ([28]). Additionally, Yang and Zheng proposed a generalization of Yau’s Conjecture 1.1 for the Hermitian case ([44, Conjecture 1.6 (b)]).
@@
We say that $(M,h)$ has positive real bisectional curvature denoted by $B_{h}>0$ if for any $x\in M$, any unitary frame $e=\{e_{1},\cdots,e_{n}\}$ at $x$ and any nonnegative constants $a=\{a_{1},\cdots,a_{n}\}$ with $|a|^{2}=a_{1}^{2}+\cdots+a_{n}^{2}>0$, it holds that
@@
where $R$ is the Chern curvature tensor of $h$. This curvature condition is equivalent to positive holomorphic sectional curvature when the metric $h$ is Kähler, and it’s slightly stronger in the non-Kähler case (cf. [44]). Under the entra assumption that $M$ is Kählerian, Conjcture 1.11 is just a weaker version of Conjecture 1.2. In this special case, Tang ([39, Theorem 1.4]) confirmed the Conjecture 1.11 for $\dim M=3$ and proved that $M$ is projective for $\dim M\geq 1$. Applying the Bochner-type formula used in the proof of Proposition 3.1, we confirm Conjecture 1.11 under the additional assumption that $M$ is Kählerian.
@@
This paper is organized as follows. In Section 2, we recall some basic definitions and facts. In Section 3, we establish an integral inequality to show Theorem 1.4 and provides the proofs of Theorem 1.5 and Theorem 1.7. In section 4, we present a proof of Theorem 1.9. In Section 5, we give a proof of Theorem 1.12. In the appendix, we provide a proof of $L^{2}$-vanishing theorem.
@@ 2. Preliminaries
For the reader’s convenience, we gather some basic definitions and useful facts relevant to the study of rational connectedness and holomorphic sectional curvature, which are well-known to experts and will be referenced later in the text
@@ 2.1. Pseudo-effective line bundle and plurisubharmonic function
Let $\Omega$ be an open subset of $\mathbb{C}^{n}$. We have the following approxiamation properties for plurisubharmonic functions (see e.g. [14, Chapter 1 and Chapter 3]).
@@
([4, Theorem 2.1]) Let $\{v_{j}^{1}\},\cdots,\{v_{j}^{k}\}$ be a decreasing sequences of functions in $Psh(\Omega)\cap L^{\infty}_{loc}(\Omega)$ and assume that for all $z\in\Omega$,
@@ 2.2. Some criterions for rational connectedness
Based on the uniruledness criterion estabished in ([7]), Campana, Demailly and Peternell devoloped the following criterion for rational connectedness.
Lemma 2.5 ([9, Theorem 1.1]).
@@ 2.3. Some properties of curvature tensor
(Royden’s observation, [38, P552]) Let $\xi_{1},\cdots,\xi_{k}$ be $k$ orthonormal tangent vectors. If $S(\xi,\bar{\eta},\zeta,\bar{\omega})$ is a symmetric bihermitian form on some complex vector space $V$ (i.e., $S(\xi,\bar{\eta},\zeta,\omega)=S(\zeta,\bar{\eta},\xi,\bar{\omega})$ and $S(\xi,\bar{\omega},\zeta,\bar{\eta})$), such that for all $\xi\in V$, we have
@@
([43, Lemma 1.6]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$. Let $e_{1}\in T_{x_{0}}^{1,0}M$ be a unit vector which minimizes the holomorphic sectional curvature $H$ of $\omega$ at $x_{0}$, then
@@
([23, Lemma 2.1]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$. Assume that the holomorphic sectional curvature $H$ of $\omega$ is nonnegative at $x_{0}$. If a nonzero tangent vector $V\in T^{1,0}_{x_{0}}M$ such that $R(V,\overline{V},W,\overline{W})=0$ for any $W\in T^{1,0}_{x_{0}}(M)$, then
@@
2.4. The structure theorem of morphism
The following theorem reveals a detailed structure of morphisms whose domain has semi-positive holomorphic sectional curvature.
Lemma 2.9.
[31, Theorem 1.6] Let $(X,g)$ be a compact Kähler manifold with semi-positive holomorphic sectional curvature, and let $\varphi:X\to Y$ be a surjective morphism to a compact Kähler manifold $Y$ with a pseudo-effective canonical bundle. Then
  1. (1)
    The following statements hold:
    1. The morphism $\varphi:X\to Y$ is smooth and locally trivial.
    2. The exact sequence of vector bundles $0\to T_{X/Y}\to TX\xrightarrow{d\varphi^{*}}\varphi^{*}TY\to 0$ admits a holomorphic orthogonal splitting, that is, the orthogonal complement $T_{X/Y}^{\perp}$ is a holomorphic vector bundle and there exists an isomorphism $j:\varphi^{*}TY\to T_{X/Y}^{\perp}$ such that it gives the holomorphic orthogonal decomposition
      $TX=T_{X/Y}\oplus j(\varphi^{*}TY)\cong T_{X/Y}\oplus\varphi^{*}TY.$
    3. The image $Y$ admits a Kähler metric $g_{Y}$ such that $g_{Q}=\varphi^{*}g_{Y}$ and the holomorphic sectional curvature of $g_{Y}$ is identically zero. Here $g_{Q}$ is the Hermitian metric on $\varphi^{*}TY$ induced by the above exact sequence and the metric $g$.
  2. (2)
    Let $F$ be a fiber of $\varphi$ and $F_{\text{univ}}$ be its universal cover. We obtain the isomorphism
    $X_{\text{univ}}\cong\mathbb{C}^{m}\times F_{\text{univ}},$
    where $m$ is the dimension of $Y$. There exists a Kähler metric $g_{F_{\text{univ}}}$ on $F_{\text{univ}}$ such that the holomorphic sectional curvature of $g_{F_{\text{univ}}}$ is semi-positive and the above isomorphism $X_{\text{univ}}\cong\mathbb{C}^{m}\times F_{\text{univ}}$ is isometric with respect to the Kähler metrics induced by $g$, $g_{Y}$, and $g_{F_{\text{univ}}}$.
3. Proof of Theorem 1.4, Theorem 1.5 and Theorem 1.7
In this section, we provide the proofs of Theorem 1.4, Theorem 1.5 and Theorem 1.7. First, we present a proof of Theorem 1.4. The proof of Theorem 1.4 is divided into two parts: establishing an integral inequality and performing a pointwise computation. The method is based on integration by parts, Bochner-type formula and algebraic properties of holomorphic sectional curvature.